# Identifying Switching of Antiferromagnets by Spin-Orbit Torques

### v1

@@ I INTRODUCTION
Antiferromagnets (AFMs) have been proposed as a groundbreaking platform in spintronics, enabling information storage through the encoding of data in the alignment of the staggered magnetization, also termed Néel vector [1, 2, 3, 4]. The integration of antiferromagnets (AFMs) as active components in spintronic devices offers significant advantages, including remarkable stability against external magnetic fields [5], and, most notably, intrinsically fast terahertz (THz) dynamics [6].
@@ II Results and discussion
The direction of NSOT driven Néel vector reorientation is fully determined by the direction of the driving current . Following the derivation by Železný et al. [7], the staggered effective fields at the AFM sublattices A and B are defined by the cross product of the normal on an easy plane and the direction of the in-plane current :
@@
The resulting canting of the sublattice magnetizations in the same direction out of the easy-plane generates an exchange torque driven rotation of the Néel vector around , whose direction depends on the canting direction [2]. is maximum for antiparallel/ parallel alignment and zero for perpendicular alignment of the Néel vector to the current direction. If the magnetization directions of the AFM sublattices are swapped (), corresponding to a sign change of the Néel vector, the torques change sign as well as shown in Eq. 2.
@@
Tetragonal Mn2Au is characterized by a strong magnetic anisotropy with a hard c-axis. The magnetic anisotropy within the (001)-plane, which is much smaller, favors two mutually orthogonal magnetic easy axes along the crystallographic directions. This results in a domain configuration of as grown epitaxial Mn2Au(001) thin film of sub-m2 sized domains with four directions of the Néel vector [19]. Based on the discussion above, an NSOT arising from a current parallel to the [110] direction is expected to switch all domains with Néel vectors aligned parallel or antiparallel to this direction, not only perpendicular to the current but also into the same direction, such as the [10] direction. As no NSOT acts on the domains with Néel vectors aligned perpendicular to the current direction, those with Néel vectors pointing in the antiparallel [10] direction are unaffected, and domain walls will persist. However, these domain walls are expected to be eliminated by a second current pulse now applied along the [10] direction, as the NSOT will again rotate the antiparallel Néel vectors in opposite directions around the thin film normal by 90o. Thus, NSOT is characterized by the complete alignment of the Néel, resulting in the formation of one large AFM domain within the active area of a switching device.
@@
In contrast, the thermomagnetoelastic mechanism results in the formation of one axis, along which the Néel vector aligns both parallel and antiparallel, i. e. it favors the formation of domains. Thermomagnetoelastic switching arises from current-induced heating and the resulting anisotropic thermal expansion of patterned thin-film samples clamped to thick insulating substrates. This can be illustrated by considering a thin, patterned metallic stripe heated by a current pulse. The normal thermal expansion of the stripe, associated with its temperature increase, is constrained by its clamping to the substrate, resulting in strain. However, this constraint is stronger in the direction parallel to the stripe than in the perpendicular direction, leading to compressive strain along the stripe. Thus, the magnetic anisotropy is modified, driving a reorientation of the Néel vector. The axis of thermomagnetoelastic Néel vector alignment depends on the specific device geometry, as demonstrate below through simulations. Consequently, it can be either parallel or perpendicular to the current direction, depending on the sample geometry, which serves as a clear criterion for distinguishing it from NSOT-driven Néel vector reorientation.
@@
The typical geometry for current induced Néel vector switching is cross-shaped. As shown in Fig. 1, this cross can be oriented with its arms parallel to the easy axes ( cross) or parallel to the in-plane hard axes ( cross). In both cases, the current direction in the center of the cross is selected to be parallel to the direction as shown by the COMSOL [20] simulations in panels a and d of Fig. 1. The simulated spatially dependent temperature distribution at the end of a current pulse with a width of 10 is shown in panels b and e. The corresponding spatial distributions of the strain, i.e. the relative difference in expansion between the and the directions are shown in panels c and f. Please note that the sign of strain is different for the two geometries, i.e. the centers of the crosses are strained (elongated) for the cross perpendicular and for the cross parallel to the current direction. Furthermore, the absolute value of the current heating induced strain is about a factor of five smaller for the cross geometry. The magnitude of the strain generally decreases with decreasing pulse widths and current density.
@@
The as-prepared AFM domain configuration of the epitaxial Mn2Au(001) thin films is discussed in detail in Ref. [21]; a typical example, featuring an equal population of sub-m2 sized domains, each with the Néel vector aligned along one of the four easy -directions, is presented in Fig. S1 of the Supplementary Information.
@@
Current pulses with a duration of 10 s, applied along the same crystallographic [10] direction, result in a strong predominance of domains with the Néel vector aligned perpendicular to the current direction (bright) for the 0o cross configuration shown in panel g. In contrast, for the 45o cross configuration shown in panel h, the Néel vector orientation parallel to the current direction (dark) is dominant. This means that the same current pulse direction leads to mutually perpendicular alignment of the Néel vector in the center of the 0o and 45o crosses, corresponding to the perpendicular directions of the current-induced strain shown in panels c and f. Thus, these results provide strong evidence for the thermomagnetoelastic switching mechanism. Moreover, our experiments rule out NSOT as the driver of the reorientation process for pulse widths of 10 s or longer (see also Supplementary Information, Fig. S2), as the direction of the NSOT depends solely on the current direction.
@@
Compelling evidence for NSOT-induced switching by nanosecond current pulses is provided by the observation of a predominantly single large domain in the center of the 45o cross, as shown in Fig. 1j. This is because NSOT drives the Néel vector to align along a specific direction removing 180o domain walls. In contrast, after a 10 s pulse, the center of the thermomagnetoelastically switched cross contains multiple domain walls with a width close to the resolution limit in the XMLD-PEEM image as shown in Fig. 1g.
To directly visualize the alignment of the Néel vector after the ns pulses, we have additionally investigated 45o crosses of Mn2Au(40 nm)/Ni80Fe20(4 nm) bilayers. As described in ref. [19], these bilayers exhibit a strong collinear exchange coupling between the Néel vector and the magnetization of the ferromagnetic Ni80Fe20. By leveraging this coupling, we indirectly determine the direction of the Néel vector using XMCD-PEEM at the Fe edge, which reveals the orientation of the Ni80Fe20 magnetization (Fig. 2, panels a and b). XMLD-PEEM at the Mn edge (panel c) confirms the assumed one-to-one correspondence between the AFM and FM domains also in the switched state.
@@
Complete NSOT switching was observed at current densities ranging from (100 ns pulses) to (2.5 ns pulses) as shown in Fig. 3, panel a. These values exceed by more than a factor of five the current density of required for thermomagnetoelastic switching with 10 s pulses. As the thermally generated strain induced by current pulses decreases with shorter pulse widths, thermomagnetoelastically driven switching eventually requires higher current densities than NSOT switching. Consistent with this behavior, for a pulse width of 100 ns, we observed NSOT switching at lower current densities, followed by thermomagnetoelastic switching at higher current densities, as indicated by the transverse resistance measurements in Fig. 3a. For 100 ns pulse widths, the induced resistance changes initially follow the behavior observed for nanosecond pulses but subsequently reverse sign. In this regime, saturation of the resistance change could not be achieved, as several samples failed under further increases in current density. However, the XMLD-PEEM image shown in Fig. 3b, captured after completing the pulse sequence depicted in panel a, clearly reveals a Néel vector alignment parallel to the current direction in the central region of the sample, consistent with expectations for thermomagnetoelastic switching.
@@
The compound Mn2Au is particularly noteworthy for its ability to exhibit both thermomagnetoelastic as well as NSOT switching within the same sample, facilitating their identification. We investigated Néel vector switching in two distinct thin-film geometries, enabling a clear separation of the two mechanisms. Specifically, in the 45o-cross geometry, thermally induced strain and the predicted NSOT act in opposition. For current pulse widths in the millisecond and microsecond ranges, we observed a directional dependence of the Néel vector reorientation that is consistent with the thermal mechanism alone. However, a transition to NSOT switching is observed for pulse widths in the nanosecond range. In this regime, the Néel vector aligns along a specific direction perpendicular to the current, which is a characteristic feature of the NSOT and could enabling novel applications in spintronics.
@@ III Methods
All experimental data shown in this manuscript were obtained investigating Mn2Au(001)(45 nm) thin films grown epitaxially on Ta(001)(13 nm)/Mo(001)(20 nm) double buffer layers on MgO(100) substrates. All layers were deposited by magnetron sputtering by the process described in detail in ref. [22]. The samples were capped with 2 nm of polycrystalline SiNx to protect them from oxidation. Optical lithography and ion beam etching were used to pattern the films into cross-shaped 50 matched coplanar wave guide structures as shown in the Supplementary Information (Fig. S3).

### v2

@@ I Introduction
Antiferromagnets (AFMs) have been proposed as a groundbreaking platform in spintronics, enabling information storage through the encoding of data in the alignment of the staggered magnetization, also termed Néel vector [1, 2, 3, 4]. The integration of antiferromagnets (AFMs) as active components in spintronic devices can offer significant advantages, such as stability against external magnetic fields requiring additionally magnetic anisotropy [5]. In particular, the magnetic domain configuration of our samples is stable in magnetic fields up to 30 T [7]. Furthermore, the intrinsically fast terahertz (THz) dynamics of AFMs [6] enables the manipulation of magnetic order by THz pulses [8].
@@
In principle, there is also a third mechanism to be considered, which is based on inhomogeneous current heating due to the anisotropic magnetoresistance (AMR) of AFM domains with different Néel vector orientation. Within this model, the domains with higher resistance, i. e. temperature, should be removed [18]. However, for the compound investigate here, this mechanism can be excluded, because the AMR of Mn2Au results in a higher resistance for a perpendicular alignment of Néel vector and current [19], which is the final configuration after the switching discussed below.
@@ II Results and discussion
We consider a thin film geometry with a magnetic hard axis perpendicular to the sample plane. The direction of NSOT driven Néel vector reorientation is fully determined by the direction of the driving current . Following the derivation by Železný et al. [9], the staggered effective fields at the AFM sublattices A and B are defined by the cross product of the normal on the easy thin film plane and the direction of the in-plane current (see Fig. 1):
@@
The resulting canting of the sublattice magnetizations in the same direction out of the easy-plane generates an exchange torque driven rotation of the Néel vector around , whose direction depends on the canting direction [2]. is maximum for antiparallel/ parallel alignment and zero for perpendicular alignment of the Néel vector to the current direction. If the magnetization directions of the AFM sublattices are swapped (), corresponding to a sign change of the Néel vector, the torques change sign as well as shown in Eq. 2. Thus, as visualized in Fig. 1, NSOT driven by a current parallel to the magnetic moments will reorient the Néel vector of -domains in the same direction perpendicular to the current.
@@
Tetragonal Mn2Au is characterized by a strong magnetic anisotropy with a hard c-axis. The magnetic anisotropy within the (001)-plane, which is much smaller, favors two mutually orthogonal magnetic easy axes along the crystallographic directions. This results in a domain configuration of as grown epitaxial Mn2Au(001) thin film of sub-m2 sized domains with four directions of the Néel vector [23]. Based on the discussion above, an NSOT arising from a current along the [0] direction is expected to switch all domains with Néel vectors aligned parallel or antiparallel to this direction, not only perpendicular to the current but also into the same direction, such as the [10] direction (Fig. 1). As no NSOT acts on the domains with Néel vectors aligned perpendicular to the current direction, those with Néel vectors pointing in the antiparallel [10] direction are unaffected, and domain walls will persist. Thus, after a single current pulse through an ensemble of domains representing all four easy -directions, the Néel vector of all, but those with [10]-orientation, will point in the [10] direction perpendicular to the current. However, these remaining domain walls are expected to be eliminated by a second current pulse now applied along the [10] direction (which is perpendicular to the direction of the first pulse), as the NSOT will again rotate the antiparallel Néel vectors in opposite directions around the thin film normal by 90o. Thus, NSOT is characterized by the complete alignment of the Néel vector, resulting in the formation of one large AFM domain within the active area of a switching device.
@@
In contrast, the thermomagnetoelastic mechanism results in the formation of one axis, along which the Néel vector aligns both parallel and antiparallel, i. e. it favors the formation of domains. Thermomagnetoelastic switching arises from current-induced heating and the resulting anisotropic thermal expansion of patterned thin-film samples clamped to thick insulating substrates. This can be illustrated by considering a thin, patterned metallic stripe heated by a current pulse. The normal thermal expansion of the stripe, associated with its temperature increase, is constrained by its clamping to the substrate, resulting in strain. However, this constraint is stronger in the direction parallel to the stripe than in the perpendicular direction, leading to compressive strain along the stripe. Although the magnitude and sign of the magnetoelastic energy of Mn2Au are unknown, it is clear that the magnetic anisotropy is modified, driving a reorientation of the Néel vector. The axis of thermomagnetoelastic Néel vector alignment depends on the specific device geometry, as demonstrate below through simulations. Consequently, it can be either parallel or perpendicular to the current direction, depending on the sample geometry, which serves as a clear criterion for distinguishing it from NSOT-driven Néel vector reorientation.
@@
The typical geometry for current induced Néel vector switching is cross-shaped. As shown in Fig. 2, this cross can be oriented with its arms parallel to the easy axes ( cross) or parallel to the in-plane hard axes ( cross). In both cases, the current direction in the center of the cross is selected to be parallel to the direction as shown by the COMSOL [24] simulations in panels a and d of Fig. 2. The simulated spatially dependent temperature distribution at the end of a current pulse with a width of 10 is shown in panels b and e. The corresponding spatial distributions of the strain, i.e. the relative difference in expansion between the and the directions are shown in panels c and f. Please note that the sign of the strain is different for the two geometries, i.e. the centers of the crosses are strained (elongated) for the cross perpendicular and for the cross parallel to the current direction. This is related to the rotated orientation of the heated zones in the two geometries. Furthermore, the absolute value of the current heating induced strain is about a factor of five smaller for the cross geometry. The magnitude of the strain generally decreases with decreasing pulse widths and current density.
The materials parameters required for the simulations (heat capacity, thermal conductivity, thermal expansion coefficient, Young’s modulus, and Poisson ratio) are unknown for Mn2Au. Thus, typical values for metals were chosen. For details, see Supplemental Material, section IV.
@@
The as-prepared AFM domain configuration of the epitaxial Mn2Au(001) thin films is discussed in detail in Ref. [25]; a typical example, featuring an equal population of sub-m2 sized domains, each with the Néel vector aligned along one of the four easy -directions, is presented in Fig. S1 of the Supplemental Material [26].
@@
Current pulses with a duration of 10 s, applied along the same crystallographic [10] direction, result in a strong predominance of domains with the Néel vector aligned perpendicular to the current direction (bright) for the 0o cross configuration shown in panel g. In contrast, for the 45o cross configuration shown in panel h, the Néel vector orientation parallel to the current direction (dark) is dominant. This means that the same current pulse direction leads to mutually perpendicular alignment of the Néel vector in the center of the 0o and 45o crosses, corresponding to the perpendicular directions of the current-induced strain shown in panels c and f. Thus, these results provide strong evidence for the thermomagnetoelastic switching mechanism. Moreover, our experiments rule out NSOT as the driver of the reorientation process for pulse widths of 10 s or longer (see also Supplemental Material, Fig. S2), as the direction of the NSOT depends solely on the current direction.
@@
Compelling evidence for NSOT-induced switching by nanosecond current pulses is provided by the observation of a predominantly single large domain in the center of the 45o cross, as shown in Fig. 2j. This is because NSOT drives the Néel vector to align along a specific direction removing 180o domain walls. In contrast, after a 10 s pulse, the center of the thermomagnetoelastically switched cross contains multiple closed loops of 180o-domain walls with a width close to the resolution limit in the XMLD-PEEM image as shown in Fig. 2g and discussed in more detail in section II of the Supplemental Material.
To directly visualize the alignment of the Néel vector after the ns pulses, we have additionally investigated 45o crosses of Mn2Au(40 nm)/Ni80Fe20(4 nm) bilayers. As described in ref. [23], these bilayers behave like an exchange spring and exhibit a strong collinear exchange coupling between the Néel vector and the magnetization of the ferromagnetic Ni80Fe20. Consequently, externally driven rotations of either vector induce a coupled rotation of the other. By leveraging this coupling, we indirectly determine the direction of the Néel vector using XMCD-PEEM at the Fe edge, which reveals the orientation of the Ni80Fe20 magnetization (Fig. 3, panels a and b). XMLD-PEEM at the Mn edge (panel c) confirms the assumed one-to-one correspondence between the AFM and FM domains also in the switched state.
@@
Complete NSOT switching was observed at current densities ranging from (100 ns pulses) to (2.5 ns pulses) as shown in Fig. 4, panel a. These values exceed by more than a factor of five the current density of required for thermomagnetoelastic switching with 10 s pulses. As the thermally generated strain induced by current pulses decreases with shorter pulse widths, thermomagnetoelastically driven switching eventually requires higher current densities than NSOT switching. Only in this cross-over regime switching driven by both mechanisms can be observed at different current densities: For a pulse width of 100 ns, we observed NSOT switching at lower current densities, followed by thermomagnetoelastic switching at higher current densities, as indicated by the transverse resistance measurements in Fig. 4a. For 100 ns pulse widths, the induced resistance changes initially follow the behavior observed for nanosecond pulses but subsequently reverse sign. In this regime, saturation of the resistance change could not be achieved, as several samples failed under further increases in current density. However, the XMLD-PEEM image shown in Fig. 4b, captured after completing the pulse sequence depicted in panel a, clearly reveals a Néel vector alignment parallel to the current direction in the central region of the sample, consistent with expectations for thermomagnetoelastic switching.
@@
The compound Mn2Au is particularly noteworthy for its ability to exhibit both thermomagnetoelastic as well as NSOT switching within the same sample, facilitating their identification. We investigated Néel vector switching in two distinct thin-film geometries, enabling a clear separation of the two mechanisms. Specifically, in the 45o-cross geometry, thermally induced strain and the predicted NSOT act in opposition. For current pulse widths in the millisecond and microsecond ranges, we observed a directional dependence of the Néel vector reorientation that is consistent with the thermal mechanism alone. However, a transition to NSOT switching is observed for pulse widths in the nanosecond range, as required for ultrafast applications. In this regime, the Néel vector aligns along a specific direction perpendicular to the current, which is a characteristic feature of the NSOT and could enabling novel applications in spintronics.
@@ III Methods
All experimental data shown in this manuscript were obtained investigating Mn2Au(001)(45 nm) thin films grown epitaxially on Ta(001)(13 nm)/Mo(001)(20 nm) double buffer layers on MgO(100) substrates. All layers were deposited by magnetron sputtering by the process described in detail in ref. [27]. The samples were capped with 2 nm of polycrystalline SiNx to protect them from oxidation. Optical lithography and ion beam etching were used to pattern the films into cross-shaped 50 matched coplanar wave guide structures as shown in the Supplemental Material (Fig. S4).